OpenStudy (anonymous):

How would you go about proving the identity of: (1/1)-[sin^2 theta/(1+cos theta)]=2 csc theta

OpenStudy (anonymous):

\[\Large \frac{1}{1}-\frac{\sin^2 \theta}{1+\cos \theta}=2 \csc \theta\]

OpenStudy (anonymous):

Is this your equation?

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \ \frac{1}{1} -\frac{\sin^2θ}{1+\cosθ} =2\cscθ }\) \(\LARGE\color{blue}{ \ \frac{1+\cosθ}{1+\cosθ} -\frac{\sin^2θ}{1+\cosθ} =2\cscθ }\) \(\LARGE\color{blue}{ \ \frac{1+\cosθ-\sin^2θ}{1+\cosθ} =2\cscθ }\)

OpenStudy (solomonzelman):

I got disconnected in the middle of typing -:(

OpenStudy (anonymous):

Yes, it is my equation.

OpenStudy (anonymous):

You're fine! I truly appreciate the help!

OpenStudy (solomonzelman):

\(\large\color{blue}{ \ 1+\cosθ-\sin^2θ =2\cscθ(1 +\cosθ)}\) I don't know how to finish, sorry.

OpenStudy (anonymous):

It's fine, you've helped me so much already! I can solve it from here :)

OpenStudy (loser66):

It's not identity, how to prove. This is counterexample x = pi/2 --> sin x = 1, sin^2 =1 and csc (x) =1 also The left hand side is \(1-\dfrac{sin^2x}{(1+cosx)}= 1-\dfrac{1}{1}=0\) while the right hand side 2csc (x) =2 If it is identity, then 0 =2????

OpenStudy (loser66):

One more question: solve or prove??

OpenStudy (anonymous):

It's listed under identities. That I was asked to verify. I'm sorry.

OpenStudy (anonymous):

It says to verify/establish

OpenStudy (loser66):

No need to say sorry, it's not your fault. Can you post the page? ( please, scan and attach)

OpenStudy (anonymous):

okay, please give me a moment.

OpenStudy (loser66):

One more thing: hehehe.. I am annoyed, right? verify or justify??

OpenStudy (anonymous):

Verify.

OpenStudy (loser66):

hihihi... I fail with you. You are not alone.

OpenStudy (anonymous):

It's problem number four.

OpenStudy (anonymous):

Thank you again.

OpenStudy (loser66):

hey, your number four has the right hand side is cos, not csc

OpenStudy (loser66):

you intermingle number3 and number 4 left hand side: number4 right hand side number 3 ha!!

OpenStudy (loser66):

so, you fail alone, me not. hihihihi

OpenStudy (anonymous):

Did I really?

OpenStudy (loser66):

check it by yourself.

OpenStudy (anonymous):

I did!! I am so sorry!

OpenStudy (loser66):

Ye, should say sorry to @SolomonZelman It's your fault.

OpenStudy (solomonzelman):

why are you saying sorry to me ? @Loser66

OpenStudy (anonymous):

I am so sorry @SolomonZelman for mistyping my problem.

OpenStudy (loser66):

me not, the asker should say sorry to you only. It's her fault, not mine. hehehe...

OpenStudy (solomonzelman):

okay... ;)